OpenStudy (anonymous):

evaluate the following expression 2(1,0)-3(2,-1) A. (-4,3) B. (2,-3) C. (2,3) D. (-3,-2)

OpenStudy (nubeer):

A

OpenStudy (anonymous):

why?

OpenStudy (nubeer):

\[2\left(\begin{matrix}1 \\ 0\end{matrix}\right) -3 \left(\begin{matrix}2 \\ -1\end{matrix}\right)\] \[\left(\begin{matrix}2 \\ 0\end{matrix}\right) - \left(\begin{matrix}6 \\ -3\end{matrix}\right)\] \[\left(\begin{matrix}2-6 \\ 0-(-3)\end{matrix}\right)\]

OpenStudy (anonymous):

oh okay I get it now Thanks

hartnn (hartnn):

explain BEFORE u give answer.

OpenStudy (nubeer):

i know.. was doing that but ty.

hartnn (hartnn):

ideally don't give out answers.